# Self-organizing Multiagent Target Enclosing under Limited Information and Safety Guarantees

### v1

@@ II-C Design of Control Objectives
that selects only one potentially colliding neighbor from the set $\mathcal{Z}_{i}$, thereby allowing us to analyze the behavior of only one pair of pursuers at any instant of time. We refer to the set $\mathcal{C}_{i}$ as the set of the nearest colliding pursuer. In essence, $\mathcal{C}_{i}\subset\mathcal{Z}_{i}\subset\mathcal{N}_{i}$.
Remark 3.
In this work, we show that if $d_{ij}>0$ for all $t\geq 0$ in addition to (4), then the pursuers will exhibit a self-organizing behavior during target enclosing while also guaranteeing safety. Ensuring $d_{ij}>0$ will imply that $r_{ij}>0$ even if $\psi_{ij}\to 0$, meaning that every pursuer will maintain a safe distance with respect to every other pursuer in its vicinity.
The scenario represented in Figure 5 considers three pursuers, $P_{i}$, $P_{j}$, and $P_{k}$ aiming to enclose the target within a circle while also ensuring safety. The dotted curves in Figure 5 represent their instantaneous loiter circles around the target. In this scenario, $P_{j}$ and $P_{k}$ initially avoid encountering a nearest colliding pursuer, thus converging to the desired proximity in the same order as they started. Meanwhile, $P_{i}$ encounters $P_{j}$ as its nearest colliding pursuer, prompting $P_{i}$ to maintain $d_{ij}>0$ to avoid collision with $P_{j}$. Subsequently, at a later time during enclosing, $\psi_{ij}\to 0$, enabling $P_{i}$ to relax its safety constraint and settle behind $P_{j}$, ultimately resulting in collision-free self-organizing target enclosing.
Figure 5: Multagent target enclosing scenario with safety constraints.
$d_{ij}>0.$ (8)
This condition guarantees that all pursuers move along a loiter circle with a radius greater than its distance from the nearest colliding pursuer around the target. As a result, $\psi_{ij}\to 0$, but $d_{ij}\not\to 0$, which implies that $r_{ij}\not\to 0$, ensuring that each pursuer keeps a safe distance from its nearest colliding pursuer and is influenced by at most one other pursuer during enclosing.
Figure 5 shows the behavior of pursuers who ensure safety conditions in (8), where $P_{j}$ is the nearest colliding pursuer for $P_{i}$. Initially, $\psi_{ij}$ starts at a negative value and becomes positive over time, which is similar to the scenario without safety conditions presented in Figure 4. However, due to the safety condition, $P_{i}$ maintains a larger distance from the target than $P_{j}$, resulting in a positive $d_{ij}$. Once $\psi_{ij}=0$, $P_{j}$ is no longer the nearest colliding pursuer for $P_{i}$, and $P_{i}$ begins to converge to the desired enclosing geometry while $\psi_{ij}$ becomes increasingly positive. At steady state, $P_{i}$ autonomously converges to a finite positive value of $\psi_{ij}$ with respect to $P_{j}$. This behavior, where pursuers converge to a circle of desired proximity from the target while maintaining a safe distance from each other, is referred to as self-organizing multiagent target enclosing behavior in our paper. Unlike previous works [8, 9, 13, 24, 25, 26], the safety condition in (8) constrains $d_{ij}$ rather than $\psi_{ij}$, allowing each pursuer to independently converge to any finite $\psi_{ij}$ with respect to its neighbors. Such behavior enables pursuers to operate without the need for a predetermined formation structure, reducing the amount of information required and coordinating efforts with just one other pursuer, which is the nearest colliding pursuer. This affords pursuers greater autonomy to maneuver independently while still maintaining a stable formation/enclosing geometry. Moreover, pursuers can enter or exit the enclosing formation seamlessly, without disturbing the overall structure.
We are now in a position to summarize the control objectives in regards to the self-organizing multiagent target enclosing under limited information and safety guarantees, that is, $\lim_{t\to\infty}\left|r_{iT}-r_{d}\right|\to 0$, and $d_{ij}>0,~{}\forall\,P_{j}\in\mathcal{C}_{i}$.
III-A Analysis of Inter-Pursuer Behavior
Figure 6: Regions around the $i$th pursuer.
Lemma 1.
Pursuers situated in Regions I and III corresponding to the $i$th pursuer’s position do not have a tendency to collide with the $i$th pursuer.
Proof.
To demonstrate this, we examine the motion of pursuers in the four regions defined in (III-A) and determine the conditions under which both $d_{ij}\to 0$ and $\psi_{ij}\to 0$ can occur. We do this by deriving the dynamics of the inter-pursuer radial distance and inter-pursuer angular spacing, which are given by,
$\dot{d}_{ij}$ $=-v\left(\cos\sigma_{i}-\cos\sigma_{j}\right),$ (10a)
obtained by differentiating (II-B) with respect to time and using (II-A). Provided that all pursuers initiate their movement away from the target and tend to move toward it to enclose the latter, we can deduce from (1a) that $\dot{r}_{iT}<0,\dot{r}_{jT}<0$. This pertains to the scenario when they are moving toward the target, thus limiting their look angles within the interval $(0,\pi/2)$. Now, if a collision is imminent, then $d_{ij}\to 0$ and $\psi_{ij}\to 0$. Let us analyze each of these conditions in the four regions to analyze the situation when every pursuer is heading toward the target to enclose it.
Using (10a), we can determine the conditions that lead to $d_{ij}\to 0$ in the four regions as follows:
Region I: $\dot{d}_{ij}<0,\text{if}\,\sigma_{i}<\sigma_{j},$
Region II: $\dot{d}_{ij}>0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region III: $\dot{d}_{ij}>0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region IV: $\dot{d}_{ij}<0,\text{if}\,\sigma_{i}<\sigma_{j}.$ (11)
Next, we analyze cases when ${\psi}_{ij}\to 0$ when the above conditions for $d_{ij}\to 0$ also hold. We denote $\sigma_{i}^{\star}$ as the look angle about which ${\psi}_{ij}$ changes its sign, given by
$\sin\sigma_{i}^{*}=\frac{r_{iT}}{r_{jT}}\sin\sigma_{j},$ (12)
obtained by equating $\dot{\psi}_{ij}$ to zero in (10b).
In Regions I and IV, we have $d_{ij}>0$ or $\dfrac{r_{iT}}{r_{jT}}>1$, which implies $\sigma_{i}^{*}>\sigma_{j}$. Thus, $\sigma_{i}<\sigma_{j}$ leads to $\sin\sigma_{i}<\dfrac{r_{iT}}{r_{jT}}\sin\sigma_{j}$, resulting in $\dot{\psi}_{ij}>0$. Similarly, in Regions II and III, we have $\sigma_{i}^{*}<\sigma_{j}$, and hence $\sigma_{i}>\sigma_{j}$ implies $\sin\sigma_{i}>\dfrac{r_{iT}}{r_{jT}}\sin\sigma_{j}$, resulting in $\dot{\psi}_{ij}<0$. Therefore, we can summarize the inferences from the current analysis and those in (III-A) as
Region I: $\dot{d}_{ij}<0,\dot{\psi}_{ij}>0,\text{if}\,\sigma_{i}<\sigma_{j},$
Region II: $\dot{d}_{ij}>0,\dot{\psi}_{ij}<0,\text{if}\,\sigma_{i}>\sigma_{j},$
It follows from (III-A) and (III-A) that inter-pursuer collision can only occur in Regions II and IV during target enclosing. Therefore, no agents in Regions I and III will collide with $P_{i}$. This concludes the proof. ∎
Lemma 1 shows that the nearby pursuers in Regions I and III will not collide with $P_{i}$, so ensuring safety with those pursuers is unnecessary for $P_{i}$. This could reduce computational complexity and make the design simpler. Now, we explain why choosing one pursuer from the surrounding pursuers according to (7) is necessary to ensure safety among all pursuers.
Theorem 1 (Inter-pursuer Safety).
Proof.
Figure 7: Engagement with neighboring pursuers in Region II of $P_{i}$.
Let us consider a situation where pursuer $P_{i}$ has another pursuer, $P_{j}$, located within its Region II (denoted as Region Ii), as shown in Figure 7. Since $P_{i}$ is the nearest colliding pursuer for $P_{j}$, according to (7), $P_{j}$ must meet the safety criterion in (8) and maneuver accordingly. Hence, $P_{i}$ can safely approach the desired proximity without colliding with pursuers in its Region IIi. This means that $P_{i}$ can focus on avoiding collisions with pursuers in its Region IVi.
Before considering the second scenario, we first analyze the inter-pursuer distance $r_{ij}$ between $P_{i}$ and $P_{j}$ to arrive at the nearest colliding pursuer selection criterion discussed in (III-A). Following (III-A), we find the values of $d_{ij}$ and $\psi_{ij}$ that minimizes $r_{ij}$ by evaluating the partial derivatives of $r_{ij}^{2}$ as in (3) with respect to $d_{ij}$ and $\psi_{ij}$. We obtain the gradient $\nabla r_{ij}^{2}$ and Hessian $\nabla^{2}r_{ij}^{2}$ as,
$\nabla r_{ij}^{2}=$ $\begin{bmatrix}\dfrac{\partial r_{ij}^{2}}{\partial d_{ij}}\\ \dfrac{\partial r_{ij}^{2}}{\partial\psi_{ij}}\end{bmatrix}=\begin{bmatrix}2d_{ij}+2r_{iT}\left(1-\cos\psi_{ij}\right)\\ 2r_{iT}^{2}\sin\psi_{ij}+2r_{iT}d_{ij}\sin\psi_{ij}\end{bmatrix},$ (18)
$\nabla^{2}r_{ij}^{2}=$ $\begin{bmatrix}2&2r_{iT}\sin\psi_{ij}\\ 2r_{iT}\sin\psi_{ij}&2r_{iT}^{2}\cos\psi_{ij}+2r_{iT}d_{ij}\cos\psi_{ij}\end{bmatrix},$ (21)
It can be readily observed from (3), (18) and (21) that the minimum value of $r_{ij}=0$ is attained only when $d_{ij}=0$ and $\psi_{ij}=0$, since at these values $\nabla r_{ij}^{2}=0$ and $\nabla^{2}r_{ij}^{2}\succ 0$ (positive definite).These inferences follow the discussions in Section II-B. From (III-A), we know that in Region IV of $P_{i}$ (Region IVi in Figure 7), $d_{ik}>0$ and $\psi_{ik}<0$ for an arbitrary pursuer $P_{k}$ located therein. Therefore, when $r_{ik}\neq 0$, choosing $P_{k}$ with the smallest $d_{ik}$ and largest $\psi_{ik}$ will select only one pursuer that is closest to $P_{i}$ in terms of inter-pursuer distance. Our approach ensures safety among pursuers and allows them to self-organize during target enclosing by first picking pursuers located on the nearest loiter circle around the target to minimize $d_{ij}$. This set of pursuers is denoted by $\mathcal{Z}_{i}$, as seen in (6). Next, we select the nearest colliding pursuer with the largest $\psi_{ij}$ from the set $\mathcal{Z}_{i}$, fulfilling the conditions in (7). A pseudocode representation of the algorithm for each pursuer to identify the nearest colliding pursuer is provided in Algorithm 1.
@@ III-A Analysis of Inter-Pursuer Behavior
In the second situation, we consider the presence of multiple pursuers ($P_{j}$, $P_{k}$ and $P_{l}$) in Region IVi, as shown in Figure 8. Given the scenario, pursuer $P_{j}$ has the shortest distance to $P_{i}$ ($r_{ij}$), while pursuer $P_{k}$ has the smallest distance between itself and another pursuer ($d_{ik}$) based on the loiter circles. Additionally, pursuer $P_{l}$ has the largest inter-pursuer angular spacing between itself and $P_{i}$ ($\psi_{il}$). This implies $d_{ik}<d_{ij}<d_{il}$, and $\psi_{il}>\psi_{ij}>\psi_{ik}$. From (7), we can infer that $P_{k}$ is the nearest colliding pursuer for $P_{i}$, who only needs to concern itself with avoiding collision with $P_{k}$ by executing a suitable maneuver.
@@
For the sake of contradiction, let us consider either $P_{j}$ or $P_{l}$ to the nearest colliding pursuer for $P_{i}$. First, considering $P_{j}$ to be the nearest colliding pursuer for $P_{i}$, we have $d_{ij}>0$ based on condition (8), which also implies $d_{il}>0$. However, this will not ensure $d_{ik}>0$ and result in a collision between $P_{i}$ and $P_{k}$. Similarly, we can show that selecting $P_{l}$ as the nearest colliding pursuer will not ensure the prevention of collision of $P_{i}$ with $P_{j}$ and $P_{k}$, since $d_{il}>0$ does not imply $d_{ij}>0$ and $d_{ik}>0$. Therefore, $P_{k}$ is the appropriate choice of nearest colliding pursuer chosen as per condition (7), which requires less information than ensuring safety with a pursuer other than the nearest colliding one (which will require additional conditions). By guaranteeing safety condition (8) with respect to $P_{k}$, we can infer that $d_{ik}>0$. This subsequently implies that $d_{ij}>0$ and $d_{il}>0$, since $P_{k}$ is closer to $P_{i}$ than $P_{j}$ and $P_{l}$. In other words, ensuring that $P_{i}$ avoids colliding with $P_{k}$ automatically ensures that $P_{i}$ avoids colliding with $P_{j}$ and $P_{l}$ as well. This concludes the proof.
Figure 8: Engagement with neighboring pursuers in Region IV of $P_{i}$.

### v2

@@ II-C Design of Control Objectives
that selects only one potentially colliding neighbor from the set $\mathcal{Z}_{i}$, thereby allowing us to analyze the behavior of only one pair of pursuers at any instant of time. We refer to the set $\mathcal{C}_{i}$ as the set of the nearest colliding pursuer. In essence, $\mathcal{C}_{i}\subset\mathcal{Z}_{i}\subset\mathcal{N}_{i}$. A pseudocode representation of the algorithm for each pursuer to identify the nearest colliding pursuer is provided below in Algorithm 1.
@@
Remark 3.
In this work, we show that if $d_{ij}>0$ for all $t\geq 0$ in addition to (4), then the pursuers will exhibit a self-organizing behavior during target enclosing while also guaranteeing safety. Ensuring $d_{ij}>0$ will imply that $r_{ij}>0$ even if $\psi_{ij}\to 0$, meaning that every pursuer will maintain a safe distance with respect to every other pursuer in its vicinity. The scenario represented in Figure 3(b) considers three pursuers, $P_{i}$, $P_{j}$, and $P_{k}$, aiming to enclose the target within a circle while also ensuring safety. The dotted curves in Figure 3(b) represent their instantaneous loiter circles around the target. In this scenario, $P_{j}$ and $P_{k}$ never encounter a nearest colliding pursuer before converging to the desired proximity, settling in the same order as they started. Meanwhile, $P_{i}$ encounters $P_{j}$ as its nearest colliding pursuer, prompting $P_{i}$ to maintain $d_{ij}>0$ to avoid collision with $P_{j}$. Subsequently, at a later time during enclosing, $\psi_{ij}\to 0$, enabling $P_{i}$ to relax its safety constraint and settle behind $P_{j}$, ultimately resulting in collision-free self-organizing target enclosing.
$d_{ij}>0.$ (8)
This condition guarantees that all pursuers move along a loiter circle with a radius greater than its distance from the nearest colliding pursuer around the target. As a result, $\psi_{ij}\to 0$, but $d_{ij}\not\to 0$, which implies that $r_{ij}\not\to 0$, ensuring that each pursuer keeps a safe distance from its nearest colliding pursuer and is influenced by at most one other pursuer during enclosing. Figure 3(b) shows the behavior of pursuers who ensure safety conditions in (8), where $P_{j}$ is the nearest colliding pursuer for $P_{i}$. Initially, $\psi_{ij}$ starts at a negative value and becomes positive over time, which is similar to the scenario without safety conditions presented in Figure 3(a). However, due to the safety condition, $P_{i}$ maintains a larger distance from the target than $P_{j}$, resulting in a positive $d_{ij}$. Once $\psi_{ij}=0$, $P_{j}$ is no longer the nearest colliding pursuer for $P_{i}$, and $P_{i}$ begins to converge to the desired enclosing geometry while $\psi_{ij}$ becomes increasingly positive. At steady state, $P_{i}$ autonomously converges to a finite positive value of $\psi_{ij}$ with respect to $P_{j}$. This behavior, where pursuers converge to a circle of desired proximity from the target while maintaining a safe distance from each other, is referred to as self-organizing multiagent target enclosing behavior in our paper. Unlike previous works [8, 9, 13, 28, 29, 33], the safety condition in (8) constrains $d_{ij}$ rather than $\psi_{ij}$, allowing each pursuer to independently converge to any finite $\psi_{ij}$ with respect to its neighbors. Such behavior enables pursuers to operate without the need for a predetermined formation structure, reducing the amount of information required and coordinating efforts with just one other pursuer, which is the nearest colliding pursuer. This affords pursuers greater autonomy to maneuver independently while still maintaining a stable formation/enclosing geometry. Moreover, pursuers can enter or exit the enclosing formation seamlessly, without disturbing the overall structure. We are now in a position to summarize the control objectives in regards to the self-organizing multiagent target enclosing under limited information and safety guarantees, that is, $\lim_{t\to\infty}\left|r_{iT}-r_{d}\right|\to 0$, and $d_{ij}>0,~{}\forall\,P_{j}\in\mathcal{C}_{i}$.
Figure 4: Regions around the $i$th pursuer.
III-A Analysis of Inter-Pursuer Behavior
Lemma 1.
Pursuers situated in Regions I and III corresponding to the $i$th pursuer’s position cannot collide with the $i$th pursuer, provided all the pursuers start either inside or outside the desired enclosing orbit.
Proof.
To demonstrate this, we examine the motion of neighbouring pursuers $P_{j}$’s in the four regions as defined in (III-A) of $P_{i}$ and determine the conditions under which both $d_{ij}\to 0$ and $\psi_{ij}\to 0$ can occur. We do this by deriving the dynamics of the inter-pursuer radial distance and inter-pursuer angular spacing, which are given by,
$\dot{d}_{ij}$ $=-v\left(\cos\sigma_{i}-\cos\sigma_{j}\right),$ (10a)
obtained by differentiating (2) with respect to time and using (1). Now, if a collision is imminent, then $d_{ij}\to 0$ and $\psi_{ij}\to 0$. Let us analyze each of these conditions in the four regions to analyze the situation when every pursuer is heading toward the target to enclose it. Using (10a), we can determine the conditions that lead to $d_{ij}\to 0$ in the four regions as follows:
Region I: $\dot{d}_{ij}<0,\text{if}\,\sigma_{i}<\sigma_{j},$
Region II: $\dot{d}_{ij}>0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region III: $\dot{d}_{ij}>0,\text{if}\,\sigma_{i}>\sigma_{j},$
Region IV: $\dot{d}_{ij}<0,\text{if}\,\sigma_{i}<\sigma_{j}.$ (11)
Next, we analyze cases when ${\psi}_{ij}\to 0$ when the above conditions for $d_{ij}\to 0$ also hold, provided that all pursuers initiate their movement outside the desired proximity and tend to move toward it to enclose the target. In such a scenario, we can deduce from (1a) that $\dot{r}_{iT}<0,\dot{r}_{jT}<0$ (because agents are moving toward the desired orbit regardless of any strategy), thus limiting their bearing angles $\sigma_{i},\sigma_{j}$ within the interval $(0,\pi/2)$. Such a consideration will facilitate easy verification of the signs of $\dot{\psi}_{ij}$ since the sine function is monotonously increasing in $(0,\pi/2)$. We denote $\sigma_{i}^{\star}$ as the look angle about which ${\psi}_{ij}$ changes its sign, given by $\sin\sigma_{i}^{*}=\frac{r_{iT}}{r_{jT}}\sin\sigma_{j},$ obtained by equating $\dot{\psi}_{ij}$ to zero in (10b). In Regions I and IV, we have $d_{ij}>0$ or $\dfrac{r_{iT}}{r_{jT}}>1$, which implies $\sigma_{i}^{*}>\sigma_{j}$. Thus, $\sigma_{i}<\sigma_{j}$ leads to $\sin\sigma_{i}<\dfrac{r_{iT}}{r_{jT}}\sin\sigma_{j}$, resulting in $\dot{\psi}_{ij}>0$. Similarly, in Regions II and III, we have $\sigma_{i}^{*}<\sigma_{j}$, and hence $\sigma_{i}>\sigma_{j}$ implies $\sin\sigma_{i}>\dfrac{r_{iT}}{r_{jT}}\sin\sigma_{j}$, resulting in $\dot{\psi}_{ij}<0$. Therefore, we can summarize the inferences from the current analysis and those in (III-A) as
Region I: $\dot{d}_{ij}<0,\dot{\psi}_{ij}>0,\text{if}\,\sigma_{i}<\sigma_{j},$
Region II: $\dot{d}_{ij}>0,\dot{\psi}_{ij}<0,\text{if}\,\sigma_{i}>\sigma_{j},$
It follows from (III-A) and (III-A) that inter-pursuer collision can only occur in Regions II and IV during target enclosing. Therefore, no agents in Regions I and III will collide with $P_{i}$. Similarly, we can analyze the signs of $\dot{\psi}_{ij}$ when all the pursuers start inside the desired proximity and move away from the target to reach the enclosing orbit to reach similar conclusions as those stated previously. Note that under such conditions, $\sigma_{i},\sigma_{j}$ must lie in the interval $(\pi/2,\pi)$ to ensure $\dot{r}_{iT}>0$ and $\dot{r}_{jT}>0$. Additionally, the sine of the bearing angles will be strictly decreasing in this interval. ∎
Remark 4.
It is worth noting that Lemma 1 does not require exact knowledge of the control law. Instead, it relies on the general behavior of the pursuers that must be exhibited by any strategy designed to enclose the target. In this work, we deliberately exclude scenarios where $P_{i}$ and its neighbor $P_{j}$ are positioned in different regions about the desired proximity (some inside and some outside the desired enclosing orbit) for simplicity and an intuitive understanding. Determining the conditions when a collision will not occur in such cases may require the exact knowledge of the control law, which we leave as a future task.
Lemma 1 shows that nearby pursuers in Regions I and III will not collide with $P_{i}$, so ensuring safety with those pursuers is unnecessary for $P_{i}$. This could make the design simpler. Now, we explain why choosing one pursuer from the surrounding pursuers according to (7) is necessary to ensure safety among all pursuers.
Theorem 1 (Inter-pursuer Safety).
Proof.
(a) Engagement with neighboring pursuers in Region II of $P_{i}$.
(b) Engagement with neighboring pursuers in Region IV of $P_{i}$.
Figure 5: Engagement in various regions.
Let us consider a situation where pursuer $P_{i}$ has another pursuer, $P_{j}$, located within its Region II (denoted as Region IIi), as shown in Figure 5(a). Since $P_{i}$ is the nearest colliding pursuer for $P_{j}$, according to (7), $P_{j}$ must meet the safety criterion in (8) and maneuver accordingly. Hence, $P_{i}$ can safely approach the desired proximity without colliding with pursuers in its Region IIi. This means that $P_{i}$ can focus on avoiding collisions with pursuers in its Region IVi.
Before considering the second scenario, we first analyze the inter-pursuer distance $r_{ij}$ between $P_{i}$ and $P_{j}$ to arrive at the nearest colliding pursuer selection criterion. To find the values of $d_{ij}$ and $\psi_{ij}$ that minimizes $r_{ij}$ by evaluating the partial derivatives of $r_{ij}^{2}$ as in (3) with respect to $d_{ij}$ and $\psi_{ij}$. We obtain the gradient $\nabla r_{ij}^{2}$ and Hessian $\nabla^{2}r_{ij}^{2}$ as, $\nabla r_{ij}^{2}=\begin{bmatrix}\dfrac{\partial r_{ij}^{2}}{\partial d_{ij}}&\dfrac{\partial r_{ij}^{2}}{\partial\psi_{ij}}\end{bmatrix}^{\top}=\begin{bmatrix}2d_{ij}+2r_{iT}\left(1-\cos\psi_{ij}\right)\\ 2r_{iT}^{2}\sin\psi_{ij}+2r_{iT}d_{ij}\sin\psi_{ij}\end{bmatrix}$, $\nabla^{2}r_{ij}^{2}=\begin{bmatrix}2&2r_{iT}\sin\psi_{ij}\\ 2r_{iT}\sin\psi_{ij}&2r_{iT}^{2}\cos\psi_{ij}+2r_{iT}d_{ij}\cos\psi_{ij}\end{bmatrix}$. It can be readily observed that the minimum value of $r_{ij}=0$ is attained only when $d_{ij}=0$ and $\psi_{ij}=0$, since at these values $\nabla r_{ij}^{2}=0$ and $\nabla^{2}r_{ij}^{2}\succ 0$ (positive definite). These inferences follow the discussions in Section II-B. Therefore, minimizing both $d_{ij}$ and $\psi_{ij}$ will single out a pursuer with minimum distance $r_{ij}$ from $P_{i}$. However, such a pursuer might not be suitable for collision avoidance mechanism ensuring $d_{ij}>0$ as in (8). Instead, our approach selects the nearest colliding pursuer as per Algorithm 1 to ensure safety among all the pursuers. We know from (III-A) that in Region IV of $P_{i}$ (Region IVi in Figure 5(a)), $d_{ik}>0$ and $\psi_{ik}<0$ for an arbitrary pursuer $P_{k}$ located therein. Therefore, when $r_{ik}\neq 0$, choosing $P_{k}$ with the smallest $d_{ik}$ and largest $\psi_{ik}$ will select only one pursuer that is closest to $P_{i}$ in terms of inter-pursuer distance. Our approach ensures safety among pursuers allowing them to self-organize during target enclosing by first picking pursuers located on the nearest loiter circle around the target to minimize $d_{ij}$. This set of pursuers is denoted by $\mathcal{Z}_{i}$, as seen in (6). Next, we select the nearest colliding pursuer with the largest $\psi_{ij}$ from the set $\mathcal{Z}_{i}$, fulfilling the conditions in (7).
@@ III-A Analysis of Inter-Pursuer Behavior
In the second situation, we consider the presence of multiple pursuers ($P_{j}$, $P_{k}$ and $P_{l}$) in Region IVi, as shown in Figure 5(b). Given the scenario, pursuer $P_{j}$ has the shortest distance to $P_{i}$ ($r_{ij}$), while pursuer $P_{k}$ has the instantaneous loiter circle closest to $P_{i}$’s instantaneous loiter circle or minimum $d_{ik}$. Additionally, pursuer $P_{l}$ has the largest inter-pursuer angular spacing between itself and $P_{i}$ ($\psi_{il}$). This implies $d_{ik}<d_{ij}<d_{il}$, and $\psi_{il}>\psi_{ij}>\psi_{ik}$. From (7), we can infer that $P_{k}$ is the nearest colliding pursuer for $P_{i}$, who only needs to concern itself with avoiding collision with $P_{k}$ by executing a suitable maneuver.
@@
For the sake of contradiction, let us consider either $P_{j}$ or $P_{l}$ to be the nearest colliding pursuer for $P_{i}$. First, considering $P_{j}$ to be the nearest colliding pursuer for $P_{i}$, we have $d_{ij}>0$ based on condition (8), which also implies $d_{il}>0$. However, this will not ensure $d_{ik}>0$ and result in a collision between $P_{i}$ and $P_{k}$. Similarly, we can show that selecting $P_{l}$ as the nearest colliding pursuer will not ensure the prevention of collision of $P_{i}$ with $P_{j}$ and $P_{k}$, since $d_{il}>0$ does not imply $d_{ij}>0$ and $d_{ik}>0$. Therefore, $P_{k}$ is the appropriate choice of pursuer chosen as per condition (7), which requires less information than ensuring safety with a pursuer other than the nearest colliding one (which will require additional conditions). By guaranteeing safety condition (8) with respect to $P_{k}$, we can infer that $d_{ik}>0$. This subsequently implies that $d_{ij}>0$ and $d_{il}>0$, since $P_{k}$ is closer to $P_{i}$ than $P_{j}$ and $P_{l}$. In other words, ensuring that $P_{i}$ avoids colliding with $P_{k}$ automatically ensures that $P_{i}$ avoids colliding with $P_{j}$ and $P_{l}$ as well. This concludes the proof. ∎